OpenStudy (anonymous):

2/3 of what number is 16?

OpenStudy (anonymous):

Let the number be X. then 2x/3 = 16

OpenStudy (anonymous):

\[16\times \frac{3}{2}\]

OpenStudy (anonymous):

2/3 x =16

OpenStudy (anonymous):

24